OpenStudy (anonymous):

Rap/hip-hop music accounted for 11.7% of the total sales of recorded music in 1998, and 13.9% of the total sales of recorded music in 2006 . Find the average rate of change of the percentage of sales of rap/hip-hop music from 1998 to 2006 . Round to the nearest hundredth of a percent. Note that the percentage sign is already given in the answer, so you do not need to enter it. The average rate of change is what percent?

OpenStudy (anonymous):

(13.9 - 11.7)/(2006 - 1998) = 2.2/8 = 0.275 = 27.5 %

OpenStudy (anonymous):

wait.. 0.275 = 0.28 percent

OpenStudy (anonymous):

let me chec

OpenStudy (anonymous):

okay great thats right. Could you maybe check out my one thts basically the sam but says rock instead of rap

OpenStudy (anonymous):

please award the medals... and post the next new question

OpenStudy (anonymous):

i wil for sure. sorry

OpenStudy (anonymous):

one sec ill get it

OpenStudy (anonymous):

Rock music accounted for 30.5% of the total sales of recorded music in 1999 , and 20.4% of the total sales of recorded music in 2009 . Find the average rate of change of the percentage of sales of rock music from 1999 to 2009 . Round to the nearest hundredth of a percent. Note that the percentage sign is already given in the answer, so you do not need to enter it. The average rate of change is .

OpenStudy (anonymous):

here it is. I will give u a medal

OpenStudy (anonymous):

please make it a new question

OpenStudy (anonymous):

dont append the second question here

OpenStudy (anonymous):

okay i will post it on the feed